OpenStudy (anonymous):

please help..whats wrong in this? let a=x a+x=2x a+x-2a=2x-2a x-a=2(x-a) 1=2...

hartnn (hartnn):

since x=a x-a =0 one cannot divide by 0

OpenStudy (anonymous):

@hartnn i think so

hartnn (hartnn):

u mean u 'thought' so ?

OpenStudy (anonymous):

nah!I said your answer makes sense

hartnn (hartnn):

ohh...yes...any more doubts ?

OpenStudy (anonymous):

just a minute...i have one

OpenStudy (anonymous):

@hartnn is 0/0=infinity?

hartnn (hartnn):

no, 0/0 is not defined.

OpenStudy (anonymous):

so... (100-100)/(100-100) is not defined 10^2-10^2/(10(10-10)) is not defined (10+10)(10-10)/(10(10-10)) is not defined ...cancelling (10-10) 10+10/10 is not defined 20/10=2 is not defined????

hartnn (hartnn):

again, cannot cancel, 10-10

OpenStudy (anonymous):

y?

hartnn (hartnn):

by cancelling 10-10=0 you are removing the part that was not defined so the part which was defined, remains, which is 2=2

OpenStudy (anonymous):

oh my god!!That thing was my doubt for 2 years...Thanks buddy...@hartnn

hartnn (hartnn):

hmmm... welcome ^_^